Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Thursday, December 10, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 10, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/10/98 (Thursday)

Segment 1 report for 12/10/98: Commissioning of the pump equipment continued at Emidio Pump Station. Trenching continued for tie-in from EMP 8.95 to 8.92. At the California Aqueduct bore pipe is being welded and x-rayed. Bore operations continued at Stream Crossing #9 (EMP 15.35). On Grapevine Rd., welding, x-ray and coating occurred from EMP's 15.51 to 15.43. Shading continued 16.11 to 16.00. On Tejon Bench, backfill continued from 16.89 to 16.64. On Lebec Rd. conduit was placed and backfill continued to 21.52. Trench was excavated to 21.58. Pipe was being pushed through the bore pit at Cuddy Creek Tributary (EMP 25.74).

Segment 2B (Castaic to Sylmar): Report for 12-10-98. Fiber optic cable was spliced near Rye Cyn Rd then tied to the electrical box serving the vault site. Near Lyons Pkwy crews backfilled the repaired section of the line with slurry. General clean up occurred at Weldon Cyn while slurry, fiber optic conduit, and coating were completed for segment 2B at EMP 71.4. Paving crews progressed and only 700' of the street remain to be paved. At Weldon Cyn crews continued to remove plate shoring and excess spoils.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 10 December 1998: 3A: NIGHT work continued at Hollywood Way. Crews were finishing grouting and placing fiber optic conduit. The crew at Tuxford Street finished pushing pipe through the bore and worked on the risers. Backfill crews continued their work between Grandview and Allen, patching up sink holes and performing general clean up. Fiber optic conduit was strung across Pacoima Wash Bridge crossing today. 3C: Crews worked north of Hubbard Avenue on the mainline skip, with welding, x-ray, and coating all working. A pull box was placed just north of Grapevine Creek. Work continued on the high speed channel cover south of Balboa Boulevard.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

Spread 1: NCRR Issued

NCRR# 1110332 Cleared NCR#110008 Date NCR Issued: 12/10/98 Monitor's Name: Etchell
Inspection Code: 3 Location:  Emidio to ANF
Begin Station: 0 End Station: 33 Landowner: Other Activity: General

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page